Solve the system by elimination. Enter you answer as an ordered pair, or type none or infinite.
-5x+3y=6
x-y=4
(I think i already posted this one but idk)

Answer by sofiyac(983)

-5x+3y=6 equation a
x-y=4 equation b
Multiply equation b by 5, leave a as is
-5x+3y=6
5x-5y=20 now add the two equations, when you do this x-terms add to 0 so you're left with y term on the left side
-2y=26 divide both sides by -2
y=-13 now plug in into either of original equations and solve for x, i'll use equation b
x-(-13)=4
x+13=4
x=-17
(-17,-13)